# The $q$-ary Gilbert–Varshamov bound can be improved for all but finitely many positive integers $q$

### 13 Mar 2024 v1

@@ 3 Hilbert class field tower and its concatenation with Lenstra codes
It was an open question in algebraic number theory, going back to Furtwängler and Hilbert, whether every number field can be embedded in a number field with class number being $1$ (i.e., equal to its Hilbert class field). In 1964, Golod and Shafarevich [11] provided a negative answer to this question by showing that the Hilbert class field tower of a number field $K$ may be infinite. An explicit example of such number field given in [11, p. 270] is a quadratic field $K=\mathbb{Q}(\sqrt{-3\cdot 5\cdot 7\cdot 11\cdot 13\cdot 17\cdot 19})=\mathbb{Q}(\sqrt{-4849845})$.
@@ 3.1 Hilbert class field tower
A main result of class field theory is that, for an algebraic number field $K$ of finite degree, there exists a maximal unramified abelian extension of $K$, denoted by $H_{K}$, in which the real places of $K$ stay real. The number field $H_{K}$ is called the Hilbert class field of $K$ (see [26, p. 399]). The most important property of $H_{K}$ is that the Galois group of the extension $H_{K}/K$ is isomorphic to the ideal class group $Cl_{K}$, i.e., $\mathrm{Gal}(H_{K}/K)\cong Cl_{K}$. There is no finite place $P\nmid\infty$ of $K$ ramified in $H_{K}$. Moreover, if $\sigma:K\hookrightarrow\mathbb{R}$ is a real embedding of $K$, then all extensions of $\sigma$ to $H_{K}$ remain real. It is clear that $H_{K}=K$ if and only if $h_{K}=1$ (i.e., $O_{K}$ is a principal ideal domain).
@@
The Golod–Shafarevich condition for the infinite Hilbert class field tower of $K$ (see Roquette [29, p. 235]) asserts that if, for some prime $p$,
@@
By Dirichlet’s unit theorem [39, p. 300], $U_{K}$ is a finitely generated abelian group and it follows that
@@ 3.2 $S_{c}$-Hilbert class field tower
Let $K$ be an algebraic number field and $S_{c}$ a finite subset of finite places of $K$, i.e., $S_{c}\subseteq S_{f}\subseteq\mathbb{P}_{K}$ with $|S_{c}|<\infty$. In the following, we present a route for theory of $S_{c}$-Hilbert class field of $K$ (see Tate [41], Cassels [5], and Neukirch [26, Chapter VI]).
@@
Before we move forward, it is mentioned that the following three group isomorphism theorems [13, p. 44] are useful in our development of theory.
@@
By the unique prime ideal factorization theorem [26, Theorem 1.3.3] and its relation with exponential valuation [26, p. 69], the group homomorphism $f:K^{*}\to I_{K}$ sending $a\in K^{*}$ to $aO_{K}\in I_{K}$ is the same as the composite of the diagonal embedding $K^{*}\hookrightarrow J_{K}$ and the group homomorphism $\pi_{f}:J_{K}\to I_{K}$ given in (22). Therefore, the image of the following composite homomorphism
@@
It follows from the finiteness of the ideal class group $Cl_{K}$ and the group isomorphism given in (29) that $C_{K}^{S}$ is a subgroup of $C_{K}$ of finite index. According to [26, Proposition 6.1.8], the subgroup $C_{K}^{S}$ of $C_{K}$ satisfies the topological requirement in the assumption of the existence theorem in the global class field theory (see, [26, Theorem 6.6.1] and [41, p. 172]). Therefore, there exists a unique finite abelian extension of $K$, denoted by $H_{K,S_{c}}$, corresponding to the subgroup $C_{K}^{S}$ of $C_{K}$. The Galois group of the abelian extension $H_{K,S_{c}}/K$ is given by
@@
By using the criteria given in [27, Proposition 2.5.3], which apply for all non-archimedean places of global fields including finite places of algebraic number fields, we see that the $S_{c}$-Hilbert class field $H_{K,S_{c}}$ of $K$ has the following two properties
@@
Let $K^{S}=K^{*}\cap J_{K}^{S}=\{a\in K^{*}\mid v_{P}(a)=0\mbox{ for all }P\in S_{f}\setminus S_{c}\}$, each element of which is called an $S$-unit of $K$. If $S_{c}=\emptyset$, i.e., $S=S_{\infty}$, then $K^{S}=K^{S_{\infty}}=K^{*}\cap J_{K}^{S_{\infty}}=U_{K}$, the group of units of $O_{K}$. The generalized Golod–Shafarevich condition has been given by Ihara [17, p. 705], which asserts that if, for some prime $p$,
@@
It was mentioned in [17, p. 705] that the proof presented in Roquette [29] for the case $S_{c}=\emptyset$ can be generalized immediately to the nonempty case $S_{c}\neq\emptyset$ just by considering the elements of $S_{c}$ as additional “infinite primes”. A proof for an even more general theorem including the generalized Golod–Shafarevich condition (30) has been briefly presented by Maire [21]. The generalized Golod–Shafarevich condition (30) for an infinite class field tower of an algebraic number field is an analogue of the one applied for an infinite class field tower of an algebraic function field over a finite field (see Schoof [32, p. 11], Niederreiter and Xing [27, p. 59], and the monograph by Serre [35]).
@@
The $p$-rank of $K^{S}$, i.e., $d_{p}K^{S}$, in (31) can be calculated as follows. According to a generalized Dirichlet’s unit theorem [26, Proposition 6.1.1], we know that $K^{S}$ is a finitely generated abelian group and that
@@
Let $K_{i}=K_{i}^{(S_{i})}$ and $\Delta_{i}=|\Delta_{K_{i}}|$ for $i\in\mathbb{N}$. Then, $[K_{i}:K]\to\infty$ as $i\to\infty$. Since each finite place $P\nmid\infty$ of $K_{i}$ is unramified in $K_{i+1}$, by virtue of an analogue of the Riemann–Hurwitz formula for algebraic number fields [26, Proposition 3.3.13], we have
@@ 5 Proof of Theorem 1
Since $q\geq Q=\exp(29)$, we have $\ell=\lfloor q^{1/6}\rfloor\geq 125$ and hence $p_{\ell}\geq p_{125}=691$. Thus, by applying an explicit bound for $\vartheta$-function given in [30, p. 211], we have
@@
6 Proof of Theorem 8
The upper bound in (8), namely $\eta(\delta)\leq 1$ for any real number $\delta\in(0,1)$, is deduced from the definition of the growth $\eta(\delta)$ in (6) and the Plotkin bound for rate function $R_{q}(\delta)$ in (5). In the following, we are going to establish the nontrivial lower bound in (8).
First, we choose
$\varepsilon_{q}=\frac{(\log q)\log\log q}{C_{0}q^{1/6}}$ (55)
$r=r(q)=\lceil(1-\varepsilon_{q})^{2}q\rceil$ (56)
for $q\geq 27$.
Next, we set
$x=\sqrt{q},\;\;\;y=(1-\varepsilon_{q})\sqrt{q}$
and choose
$\ell=\ell(q)=\lfloor x^{1/3}\rfloor$ (57)
for $q\geq 27$.
Finally, we define
for $q\in\mathbb{N}$ such that $\ell(q)\geq 10$.
We are going to prove that the following cardinality of a subset
satisfies $N_{q}\geq 2k$ for $q$ large enough.
for $q$ large enough.
Therefore, we have
Hence, for $q$ sufficiently large,
$N_{q}\geq\pi(x;4,3)-\pi(y;4,3)$ (59)
where $x=\sqrt{q}$ and $y=(1-\varepsilon_{q})x$.
$h=\varepsilon_{q}x\sim C_{1}x^{2/3}(\log x)\log\log x$ (61)
for $q$ sufficiently large with a suitable constant $C_{1}>0$.
By using theory on the distribution of primes in arithmetic progressions and in short intervals (see, e.g., [2, Theorem 3], [25, p. 656, Theorem]), we know that, for any fixed real number $C_{2}\in(0,1)$,
It follows from (59), (62), and (61) that
$N_{q}\geq C_{2}\frac{h}{2\log x}\sim C_{3}x^{2/3}\log\log x$
$2k\sim\frac{1}{2}\ell^{2}\sim\frac{1}{2}x^{2/3}$
for $q$ sufficiently large. Hence, $N_{q}\geq 2k$ for $q$ large enough.
Now, by using (38), we obtain
for $q\geq C_{4}$ and $10<\delta<$ 1, where $C_{4}>0$ is an absolute constant.
It is clear from (56) that
for $q$ sufficiently large.
Moreover, we can see that
as $q\to\infty$, where $C_{5}>0$ is some constant.
Hence, for $q$ large enough,
$\frac{\log D}{2k}<2C_{0}\varepsilon_{q}.$ (65)
Consequently, there exists an absolute constant $C>5$ such that
$R_{q}(\delta)>1-\delta-C\frac{\log\log q}{q^{1/6}}$
for all $q\geq 27$ and $10<\delta<$ 1. Hence,
$\eta(\delta)=\liminf_{q\rightarrow\infty}\frac{1}{\log q}\log[1-\delta-R_{q}(\delta)]^{-1}\geq\frac{1}{6}$
for $10<\delta<$ 1.
Therefore, the proof of Theorem 8 is complete.  $\Box$

### 21 Mar 2024 v2

@@ 3 Hilbert class field tower and its concatenation with Lenstra codes
It was an open question in algebraic number theory, going back to Furtwängler and Hilbert, whether every number field can be embedded in a number field with class number being $1$ (i.e., equal to its Hilbert class field). In 1964, Golod and Shafarevich [14] provided a negative answer to this question by showing that the Hilbert class field tower of a number field $K$ may be infinite. An explicit example of such number field given in [14, p. 270] is a quadratic field $K=\mathbb{Q}(\sqrt{-3\cdot 5\cdot 7\cdot 11\cdot 13\cdot 17\cdot 19})=\mathbb{Q}(\sqrt{-4849845})$.
@@ 3.1 Hilbert class field tower
A main result of class field theory is that, for an algebraic number field $K$ of finite degree, there exists a maximal unramified abelian extension of $K$, denoted by $H_{K}$, in which the real places of $K$ stay real. The number field $H_{K}$ is called the Hilbert class field of $K$ (see [31, p. 399]). The most important property of $H_{K}$ is that the Galois group of the extension $H_{K}/K$ is isomorphic to the ideal class group $Cl_{K}$, i.e., $\mathrm{Gal}(H_{K}/K)\cong Cl_{K}$. There is no finite place $P\nmid\infty$ of $K$ ramified in $H_{K}$. Moreover, if $\sigma:K\hookrightarrow\mathbb{R}$ is a real embedding of $K$, then all extensions of $\sigma$ to $H_{K}$ remain real. It is clear that $H_{K}=K$ if and only if $h_{K}=1$ (i.e., $O_{K}$ is a principal ideal domain).
@@
The Golod–Shafarevich condition for the infinite Hilbert class field tower of $K$ (see Roquette [34, p. 235]) asserts that if, for some prime $p$,
@@
By Dirichlet’s unit theorem [44, p. 300], $U_{K}$ is a finitely generated abelian group and it follows that
@@ 3.2 $S_{c}$-Hilbert class field tower
Let $K$ be an algebraic number field and $S_{c}$ a finite subset of finite places of $K$, i.e., $S_{c}\subseteq S_{f}\subseteq\mathbb{P}_{K}$ with $|S_{c}|<\infty$. In the following, we present a route for theory of $S_{c}$-Hilbert class field of $K$ (see Tate [46], Cassels [5], and Neukirch [31, Chapter VI]).
@@
Before we move forward, it is mentioned that the following three group isomorphism theorems [16, p. 44] are useful in our development of theory.
@@
By the unique prime ideal factorization theorem [31, Theorem 1.3.3] and its relation with exponential valuation [31, p. 69], the group homomorphism $f:K^{*}\to I_{K}$ sending $a\in K^{*}$ to $aO_{K}\in I_{K}$ is the same as the composite of the diagonal embedding $K^{*}\hookrightarrow J_{K}$ and the group homomorphism $\pi_{f}:J_{K}\to I_{K}$ given in (23). Therefore, the image of the following composite homomorphism
@@
It follows from the finiteness of the ideal class group $Cl_{K}$ and the group isomorphism given in (30) that $C_{K}^{S}$ is a subgroup of $C_{K}$ of finite index. According to [31, Proposition 6.1.8], the subgroup $C_{K}^{S}$ of $C_{K}$ satisfies the topological requirement in the assumption of the existence theorem in the global class field theory (see, [31, Theorem 6.6.1] and [46, p. 172]). Therefore, there exists a unique finite abelian extension of $K$, denoted by $H_{K,S_{c}}$, corresponding to the subgroup $C_{K}^{S}$ of $C_{K}$. The Galois group of the abelian extension $H_{K,S_{c}}/K$ is given by
@@
By using the criteria given in [32, Proposition 2.5.3], which apply for all non-archimedean places of global fields including finite places of algebraic number fields, we see that the $S_{c}$-Hilbert class field $H_{K,S_{c}}$ of $K$ has the following two properties
@@
Let $K^{S}=K^{*}\cap J_{K}^{S}=\{a\in K^{*}\mid v_{P}(a)=0\mbox{ for all }P\in S_{f}\setminus S_{c}\}$, each element of which is called an $S$-unit of $K$. If $S_{c}=\emptyset$, i.e., $S=S_{\infty}$, then $K^{S}=K^{S_{\infty}}=K^{*}\cap J_{K}^{S_{\infty}}=U_{K}$, the group of units of $O_{K}$. The generalized Golod–Shafarevich condition has been given by Ihara [20, p. 705], which asserts that if, for some prime $p$,
@@
It was mentioned in [20, p. 705] that the proof presented in Roquette [34] for the case $S_{c}=\emptyset$ can be generalized immediately to the nonempty case $S_{c}\neq\emptyset$ just by considering the elements of $S_{c}$ as additional “infinite primes”. A proof for an even more general theorem including the generalized Golod–Shafarevich condition (31) has been briefly presented by Maire [25]. The generalized Golod–Shafarevich condition (31) for an infinite class field tower of an algebraic number field is an analogue of the one applied for an infinite class field tower of an algebraic function field over a finite field (see Schoof [37, p. 11], Niederreiter and Xing [32, p. 59], and the monograph by Serre [40]).
@@
The $p$-rank of $K^{S}$, i.e., $d_{p}K^{S}$, in (32) can be calculated as follows. According to a generalized Dirichlet’s unit theorem [31, Proposition 6.1.1], we know that $K^{S}$ is a finitely generated abelian group and that
@@
Let $K_{i}=K_{i}^{(S_{i})}$ and $\Delta_{i}=|\Delta_{K_{i}}|$ for $i\in\mathbb{N}$. Then, $[K_{i}:K]\to\infty$ as $i\to\infty$. Since each finite place $P\nmid\infty$ of $K_{i}$ is unramified in $K_{i+1}$, by virtue of an analogue of the Riemann–Hurwitz formula for algebraic number fields [31, Proposition 3.3.13], we have
@@ 4 An applicable lower bound for the rate function
5 Proof of Theorem 1
First, we choose
$\varepsilon_{q}=\frac{(\log q)\log\log q}{C_{0}q^{1/6}}$ (42)
$r=r(q)=\lceil(1-\varepsilon_{q})^{2}q\rceil$ (43)
for $q\geq 27$.
Next, we set
$x=\sqrt{q},\;\;\;y=(1-\varepsilon_{q})\sqrt{q}$
and choose
$\ell=\ell(q)=\lfloor x^{1/3}\rfloor$ (44)
for $q\geq 27$.
Finally, we define
for $q\in\mathbb{N}$ such that $\ell(q)\geq 10$.
Let $\pi(x;4,3)$ denote the number of primes less than or equal to $x$ which are congruent to $3\bmod 4$.
We are going to prove that the following cardinality of a subset
satisfies $N_{q}\geq 2k$ for $q$ large enough.
for $q$ large enough.
Therefore, we have
Hence, for $q$ sufficiently large,
$N_{q}\geq\pi(x;4,3)-\pi(y;4,3)$ (46)
where $x=\sqrt{q}$ and $y=(1-\varepsilon_{q})x$.
$h=\varepsilon_{q}x\sim C_{1}x^{2/3}(\log x)\log\log x$ (48)
for $q$ sufficiently large with a suitable constant $C_{1}>0$.
By using theory on the distribution of primes in arithmetic progressions and in short intervals (see, e.g., [2, Theorem 3], [30, p. 656, Theorem]), we know that, for any fixed real number $C_{2}\in(0,1)$,
It follows from (46), (49), and (48) that
$N_{q}\geq C_{2}\frac{h}{2\log x}\sim C_{3}x^{2/3}\log\log x$
$2k\sim\frac{1}{2}\ell^{2}\sim\frac{1}{2}x^{2/3}$
for $q$ sufficiently large. Hence, $N_{q}\geq 2k$ for $q$ large enough.
Now, by using (39), we obtain
for $q\geq C_{4}$ and $10<\delta<$ 1, where $C_{4}>0$ is an absolute constant.
It is clear from (43) that
for $q$ sufficiently large.
Moreover, we can see that
as $q\to\infty$, where $C_{5}>0$ is some constant.
Hence, for $q$ large enough,
$\frac{\log D}{2k}<2C_{0}\varepsilon_{q}.$ (52)
Consequently, there exists an absolute constant $C>5$ such that
$R_{q}(\delta)>1-\delta-C\frac{\log\log q}{q^{1/6}}$
for all $q\geq 3$ and $10<\delta<$ 1.
Therefore, the proof of Theorem 1 is complete.  $\Box$
Remark 1
The estimate of (49) can be proved directly as a straightforward generalization of Ingham’s proof in [21]. In this scenario, the Dirichlet $L$-function $L(s,\chi_{4})$ of $s\in\mathbb{C}$ will be involved, where $\chi_{4}$ is the primitive Dirichlet character modulo 4. It is noted that $L(s,\chi_{4})$ has no Landau–Siegel zero. Then, the explict formula (see, e.g., [9]) and the zero-density estimates for $L(s,\chi_{4})$ (see, e.g., [28]) will enable us to prove (49). In this way, the absolute constant $C_{0}>0$ for sufficiently large $x\geq C_{0}$ satisfying (49) can be effectively computed.
@@ 6 Proof of Theorem 7
Since $q\geq Q=\exp(29)$, we have $\ell=\lfloor q^{1/6}\rfloor\geq 125$ and hence $p_{\ell}\geq p_{125}=691$. Thus, by applying an explicit bound for $\vartheta$-function given in [35, p. 211], we have
@@
7 Acknowledgment
The author would like to thank Alexander Barg for many useful conversations and suggestions. Specifically, the author is grateful for his pointing out the reference [53] and providing the author some guidance in understanding the result therein.